### v1

# Push and Pull: A Framework for Measuring Attentional Agency

Abstract.
We propose a framework for measuring attentional agency—the ability to allocate one’s attention according to personal desires, goals, and intentions—on digital platforms. Platforms extend people’s limited powers of attention by extrapolating their preferences to large collections of previously unconsidered informational objects. However, platforms typically also allow people to influence one another’s attention. We introduce a formal framework for measuring how much a given platform empowers people to both pull information into their own attentional field and push information into the attentional fields of others. We also use these definitions to shed light on the implications of generative foundation models, which enable users to bypass the implicit “attentional bargain” that underlies embedded advertising and other methods for capturing economic value from informational goods. We conclude with a set of policy strategies that can be used to understand and reshape the distribution of attentional agency online.
1. Introduction
“The summation of human experience is being expanded at a prodigious rate, and the means we use for threading through the consequent maze to the momentarily important item is the same as was used in the days of square-rigged ships… Consider a future device… in which an individual stores all his books, records, and communications, and which is mechanized so that it may be consulted with exceeding speed and flexibility.”
—Vannevar Bush, As We May Think, 1945
“In an information-rich world, the wealth of information means a dearth of something else: a scarcity of whatever it is that information consumes. What information consumes is rather obvious: it consumes the attention of its recipients. Hence a wealth of information creates a poverty of attention and the need to allocate that attention efficiently…”
—Herbert Simon, Designing Organizations for an Information-Rich World, 1971
The original promise of the web was to develop a vast interconnected network of information that could, in the words of Tim Burners-Lee, “grow and evolve with the organization and the projects it describes” (Berners-Lee, 1989). Early proponents of expanding the web “world wide” via the internet presumed that knowledge was the primary bottleneck constraining technological ingenuity, economic development, and the spread of democratic values. Collect together and make accessible all the world’s informational resources, the thinking went, and there would be no limit to what humanity could accomplish.
There are a few fundamental challenges with implementing this vision. First, as Herbert Simon (Simon, 1971) so eloquently pointed out, the limitations imposed by human attention constrain the total amount of information that people can actually make use of at any given moment, meaning that the methods by which information is organized, sorted, indexed, and navigated play a central role in determining its practical utility. Indeed, the limitations imposed by human attention mean that adding new information to the internet without concomitant advances in search can actually reduce its overall functionality. Past a certain scale, attentional limitations necessitate curation and, inevitably, grant influence to whoever controls the process by which information is searched. We call this the internet’s attention problem.
The second major complication with the kind of informational web original laid out by Tim Berners-Lee (Berners-Lee, 1989) is that it constitutes what economists call a public good: a resource that is both non-excludable (anyone with access to the internet can benefit from the public information it contains) and non-rivalrous (one person “consuming” information does not reduce its availability to be “consumed” by another person). Anyone who invests personal toil and trouble into creating and maintaining a high-quality content benefits all of humanity, but only receives an economic benefit if they find some way to monetize their contribution—web traffic, on its own, does not pay the bills. In some cases, users will provide content without extrinsic incentives (to satisfy a variety of intrinsic motivations (Redi et al., 2020)). In general, however, public goods tend to be under-provided relative to their socially optimal levels “at a price of zero.” This problem applies both to the creation of valuable content and the provision of technologies that enable one to sort and filter that content. We call this the internet’s incentive problem.
In theory, the internet is a neutral suite of protocols that enable people to freely transmit information between computers on demand. In practice, however, the internet is defined by a collection of platforms that have been built on top of these protocols and that mediate which information flows actually take place. A key insight motivating the present work is that many prominent platform categories (e.g., search engines, content libraries, social media sites, matching platforms, online marketplaces, and, increasingly, consumer AI systems) create value by helping people overcome the attention and incentive problems outlined above. Not only this, but many platforms accomplish this a very similar way—namely, by inferring people’s wants and desires, then extrapolating those preferences to intractably large sets of unconsidered informational objects (e.g., websites, products, people, content, messages, stories, and posts) (Resnick and Varian, 1997; Shapiro and Varian, 1999). To echo Vannevar Bush, platforms have become the “means we use for threading through the consequent maze to the momentarily important item” (Bush et al., 1945).
The choices platforms make about how to infer and extrapolate people’s preferences play a pivotal role in determining who has attentional agency on the internet and, increasingly, the world at large. Informally, we define attentional agency as the ability to distribute attention according to one’s own desires, goals, and intentions. In what follows, we introduce a simple mathematical framework that formalizes and operationalizes this concept. Our framework: (1) clarifies relevant aspects of how various institutions currently solve the attention and incentive problems on the internet; (2) furnishes operational measures of attentional agency that can be used to identify disparities in informational agency across distinct groups of users for any given platform, and (3) lends insight into the implications of generative foundation models and other advances in artificial intelligence.
The fact that many prominent platforms are oriented around solving the same underlying preference inference and extrapolation problem enables us to formalize attentional agency using a common mathematical framework—a generalization of the “discounted cumulative gain” formula that defines the canonical ranking problem and is used in practice as a training objective (Järvelin and Kekäläinen, 2002). The general nature of this approach enables us to abstract away from many idiosyncrasies that distinguish platforms (e.g., whether they sort websites, videos, or restaurants) and cut through to essential questions of equity and access on the modern internet: who has attentional agency on a platform, and why?
Our analysis begins with the observation that many information flows on the internet take the form of a stylized interaction between three parties:
  1. (1)
    An agent seeking to allocate their limited attention to informational objects that they themselves value highly. Prototypical examples include a user searching the web, browsing a content library, or scrolling through a social media feed.
  2. (2)
    An advocate seeking to influence what enters the agent’s limited attentional field. Prototypical examples include an advertiser, influencer, or government agency issuing a public-service announcement.
  3. (3)
    A platform, which seeks to provide value to both agents (by inferring and extrapolating their preferences to large collections of mostly unconsidered informational objects) and advocates (by allowing them to influence what agents ultimately attend to). Prototypical examples include a search engine, content library, online marketplace, social media platform, and matching platform.
On some platforms, agents and advocates are non-overlapping groups of users (e.g., job-seekers versus employers on a digital marketplace); on others, however, they are the same users acting in different capacities, either simultaneously or at different points in time (e.g., users of a dating app who both view profiles and want their own profile to be widely viewed).
When an agent and advocate have sharply conflicting priorities, a platform cannot deliver maximal value to both simultaneously. An influencer trying to “make a name for themselves,” for example, might want their work to be seen by as many people as possible, while users only want to see content that is narrowly related to their personal interests. We define two separate measures that capture how much attentional value a platform delivers to both agents and advocates.
  1. Pull:
    The amount of attentional value a platform delivers an agent as a fraction of what it is technologically prepared to deliver them given its current knowledge of agent preferences.
  2. Push:
    The amount of attentional value a platform delivers an advocate as a fraction of what it is technologically prepared to deliver them given its current knowledge of agent preferences.
Push and pull are not necessarily zero sum. Indeed, many forms of social interactions allow people to unilaterally enter into the attentional fields of others (e.g., the practice of sending a “cold-email,” which can have net positive value for the recipient). In practice, however, increasing push often does reduce pull on digital platforms: limited attention means that pushing information into an agent’s attentional field tends to exclude other things they would have preferred to see more. As we elaborate in Section 3.2, the relative alignment of agent and advocate preferences and other primitives of a platform environment define the rate at which these two measures trade off against one another.
Our framework defines pull or push at the level of the individual agent or advocate. These estimates can then be used to compare the experience of specific sub-populations as defined by age, race, gender, and other characteristics. Importantly, we define our measures in terms of representations that platforms already compute in the process of mediating informational pushes and pulls—they are therefore not only operationalizable, but in a meaningful sense already operationalized.
Our framework highlights two separate channels by which generative foundation models and other advances in artificial intelligence impact the nature of attentional agency. First, greater intelligence enables platforms to ascertain progressively more granular representation of people’s preferences and intentions leading up to instances of information retrieval. From the perspective of attentional agency, this is potentially a double-edged sword—more accurate representations of a user’s type not only enables platforms to deliver more targeted informational pulls, but also more targeted informational pushes. Our framework reveals that the ultimate agency implications of these innovations depends on the interaction of a variety of factors, including the alignment of intentions between those seeking to push and pull information and the relative accuracy with which platforms can ascertain informational relevance for each.
Second, generative foundation models excel at inferring the significance of complex digital objects at the level of their atomic informational components (“tokens”) and reconstituting them in novel ways (Bommasani et al., 2021; Bubeck et al., 2023; Yin et al., 2023). Generative foundation models therefore enable platforms to source, extract, and recombine information more flexibly. For example, whereas traditional search engines ranked relevant web pages in their entirety, AI-powered search tools can pull, rank, sort, and recombine information from within a variety of different pages on demand. This poses a fundamental challenge to the existing social, economic, and legal institutions that have emerged to regulate the capture and distribution of value on the internet. Most notably, the ability of generative foundation models to surgically extract information from its original delivery format—thus leaving embedded advertising, branding, and other promotional material behind—undermines many existing methods that platforms and content providers rely on to monetize their services. This is especially disruptive of the advertising model, which relies on the fact that people apprehend traditional informational bundles (e.g., web pages) in their entirety. Stripping away context can also undermines established norms and expectations about how information should be interpreted, modified, and shared (Jain et al., 2023).
The rest of the paper is organized as follows. In Section 2, we first review related literature. Section 3 then presents our framework and key definitions. Section 4.1 highlights the broad applicability of our framework by detailing how five major categories of current internet platforms—search, content, social media, matching, and marketplaces—map onto the primitives of our model. We then turn to the question of how generative foundation models will impact attentional agency on the internet in Section 4.2. Section 5 presents some of the immediate policy implications of the concepts and formal definitions we propose and highlights open research questions.
2. Related Work
2.1. Algorithmic Harms and Fairness
The potential harms of digital platforms have become a central topic in modern computer science (see Shelby et al. for an overview and taxonomy (Shelby et al., 2023)). Similar to our goal of operationalizing attentional agency on digital platforms, much of this work seek to translate notions such as broad beneficence, equity, and fairness into new technological domains to ensure that they are upheld (Kazim and Koshiyama, 2021).
For example, Benthall and Shekman (Benthall and Shekman, 2023) propose to extend legal fiduciary duties to the design of artificial intelligence systems by requiring that they both accurately learn people’s preferences and respect concepts such as loyalty and care. Bar-Gill et al. argue that algorithms can accelerate the capacity of firms and other sophisticated actors to exploit identifiable groups (Bar-Gill et al., 2023). Similarly, Rakova et al. propose to introduce “meaningful meaningful consent and contestability” into the legal and social relationship between online platforms and users.
Our measures of attentional agency support these existing approaches to aligning platforms to user preferences. The general perspective we develop also dominance of information ranking on the modern internet, which highlights the relevance of theoretical work on fair ranking (e.g., (Hafalir et al., 2021; Patro et al., 2022)) for broader issues of equity and access on the internet.
2.2. Economic, Psychological, and Social Value of Attention
People’s preferences over what they pay attention to arise from a wide variety of psychological, social, and economic considerations. To take but a few examples:
  • Emotional: What we pay attention directly affects our well-being through hedonic states such as curiosity, boredom, flow, mental effort, disgust, outrage, anxiety, excitement, jealousy, and regret (Loewenstein and Wojtowicz, 2023).
  • Cognitive: Paying attention to what others attend to (“joint attention”) is a deep psychological primitive that drives childhood development and adult social interactions (Mundy and Newell, 2007; Frischen et al., 2007). Simultaneously attending to common objects in physical or digital space is a key part of what makes our shared reality shared as such (Tomasello et al., 2005; Chwe, 2013).
  • Social: Receiving positive attention from others is both socially advantageous and psychologically gratifying (Braudy, 1997; Meshi et al., 2013; Lindström et al., 2021). Conversely, being ostracized or publicly shamed are among the most painful experiences one can endure (Zadro et al., 2004; Dickerson et al., 2004). The “fear of missing out” and related emotions, in particular, seem to play a key role in motivating social media use (Bursztyn et al., 2023; Abel et al., 2016).
  • Decisional: Attention is necessary to both absorb and process decision-relevant information (Stigler, 1961).
  • Economic: One’s attention can be rented to firms (the basis of many “white collar” jobs).
  • Educational: Attention is necessary to learn and make other investments in one’s human capital, especially through formal education.
Our framework distills these diverse motivations into to a single utility function and asks: to what degree to platforms maximize people’s attentional priorities given their current information and technology?
By operationalizing attentional agency in terms inferred utility values, our framework inherits both the strengths and weaknesses of revealed preference analysis. On the one hand, this approach enables us to capture a wide variety of motivations in a single measure that can be inferred from people’s behavior using standard techniques (indeed, the very techniques that platforms already use to deliver value to users). On the other hand, our measures leave open the question of how tightly coupled people’s attentional preferences are to their personal well-being or broader notions of social welfare.
Indeed, there are many situations where the two seem to diverge quite sharply: attentional addiction; highly engaging but politically divisive content; people seeking information with the intention of harming themselves or others; etc. Although our framework furnishes an operational measure of attentional agency, the question of how to evaluate its desirability in various circumstances is nuanced and requires careful consideration. Some of these concerns are only becoming more acute as AI search and other innovations enable platforms to deliver narrowly targeted persuasive messaging (Durmus et al., 2024; Salvi et al., 2024) and other novel forms of harmful content to users—an issue we return to in Section 5.2.
2.3. Information Retrieval
Our approach is motivated by the observation that information retrieval plays a central role in modern internet platforms (see Table 1). Solutions to this problem follow a common approach: (1) assign user-specific “relevance” scores to various pieces of informational content, (2) use those scores to sort content; (3) arrange items in an interface in such a way as to preferentially guide attention to items with higher rankings. Formal methods for generating and evaluating the quality of ranked content lists, known as “recommender systems,” have long been a topic of study in theoretical and applied computer science (Herlocker et al., 2004).
Approaches to recommendation include measuring the precision and recall of a list of a fixed size (e.g., the “ten blue links” returned by a search engine), considering the mean reciprocal rank (applicable to social media feeds of arbitrary extent), and variants of the discounted cumulative gain, which calculate a weighted average of relevance over ranks. Researchers have also explored methods for understanding how allocation impacts item diversity, bias, and other features of interest (Parapar and Radlinski, 2021; McAuley, 2022). The work of Ref. (Stray et al., 2021), in particular, points out that recommender systems are a special case of the more general value alignment problem and emphasizes the importance of developing operational measures of alignment—the central task of this paper.
Although the exact correspondence between offline ranking metrics and company revenues is typically proprietary, some researchers have sought to describe how offline metrics map to business value, ultimately concluding (as we might expect) that whether a given ranking metric works well is context-dependent (Jannach and Jugovac, 2019). Published work from Netflix has also discussed the role of offline experimentation and the use of offline ranking metrics like NDCG to predict business value (Alvino and Basilico, 2015; Gomez-Uribe and Hunt, 2016).
3. Formal Framework
3.1. General Setup and Key Definitions
We study a canonical interaction between three parties: an agent, who wants to pay attention to informational objects they themselves value highly; an advocate, who also has preferences about what the agent attends to; and a platform that determines whether and how informational objects enter the agent’s attentional field.
The agent has type $\theta$ drawn from a distribution $F\in\Delta(\Theta)$ where $\Theta$ is a finite type space. There is a universe of informational objects represented by an ordered set $X=(x_{1},x_{2},\dots)$. An allocation is a bijective function $\phi:\mathbb{N}\to X$. The agent has a type-dependent value for each informational object that is represented by a utility function $u:X\times\Theta\to\mathbb{R}_{\geq 0}$. Owing to their limited attention, the agent discounts informational objects based on how they are allocated according to a weakly decreasing function $\delta:\mathbb{N}\to[0,1]$ where $\delta_{0}=1$ without loss of generality. The agent’s value of allocation $\phi$ is
(1) $U(\phi,\theta)=\sum_{n\in\mathbb{N}}\delta_{n}\,u(\phi_{n},\theta)$
Logarithmic functions such as $\delta_{n}=\log_{2}(n+1)^{-1}$ have been studied extensively in the literature on “discounted cumulative gain” and are particularly influential in the design of ranking algorithms (Järvelin and Kekäläinen, 2002). More generally, different specifications of $\delta$ enable the model to capture the reduced-form effects of limited attention in a variety of problem domains and platform interfaces. For example, discounted cumulative gain is natural in situations where the agent sequentially attends to informational objects (e.g., a web search or social media feed); however, a “cutoff” function where $\delta_{n}=1$ for $n\leq N$ and $\delta_{n}=0$ otherwise might be more natural for an interface that simultaneously presented a handful of recommendations (e.g., a homepage featuring $N$ personalized suggestions, or an AI-augmented search result that synthesizes information into a block of text, all of which is presumed to be read).
The platform supplies the agent with an allocation. Let $\Gamma$ denote a partition of $X$, which induces what we will call a allocation technology. [^1] A partition of $X$ is a set of subsets of $X$ (“partitions” or “blocks”) such that every element of $X$ belong to one and only one block. Specifically, we assume that the platform can produce allocations that rearrange the order of the partition blocks of $\Gamma$, but not move objects within blocks. The interpretation is that digital information is naturally bundled together into objects (e.g., web pages, academic articles, social media posts), and greater degrees of technological sophistication are required to break apart, comprehend, extrapolate preferences to, and recombine more granular units of information within these objects. Let $\Phi(\Gamma)=\{\phi:\mathbb{N}\to X\,|\,\forall\gamma\in\Gamma,\text{ if }x_{i},x_{j}\in\gamma\text{ then }\phi^{-1}(x_{i})-\phi^{-1}(x_{j})=i-j\}$ denote the set of all possible allocations that can be constructed by permuting the blocks of $\Gamma$.
The platform has an imperfect understanding of $\theta$ and, therefore, the agent’s preferences over allocations. Let $m$ denote the platform’s model of the agent, which is a reduced-form representation of: (1) all information the platform possesses that pertains to the agent’s type (e.g., as collected from past interactions, present context, and their query); and (2) whatever algorithms and other technology the platform uses to transform this information into a prediction about their type. Some systems enable users to explicitly reveal aspects of their stable or transient preferences (e.g., by selecting interests when first creating one’s profile or entering a search query), whereas other systems implicitly learn a user’s type from patterns of engagement (e.g., dwell time or likes in a feed-based interface).
The advocate also has preferences about what the agent sees, which we represent with a second utility function $v:X\times\Theta\to\mathbb{R}_{\geq 0}$. Paralleling our notation for the agent, the advocate receives value
(2) $V(\phi,\theta)=\sum_{n\in\mathbb{N}}\delta_{n}\,v(\phi_{n},\theta)$
Note that we are assuming the advocate discounts utility using the same positional function as the agent. This is natural, for example, if the discount function is taken to represent the amount of attention the agent pays to each position and both parties value informational objects based on how likely they are to be attended to. We assume the platform chooses an allocation that maximizes a linear combination of agent and advocate utility subject to technological feasibility.
(3) $\phi_{\lambda}\in\operatorname*{arg\,max}_{\phi\in\Phi(\Gamma)}\quad\mathbb{E}\big{[}\lambda U(\phi,\theta)+(1-\lambda)V(\phi,\theta)\,\big{|}\,m\big{]}$
for a given weight $\lambda\in[0,1]$. For simplicity, we will assume throughout that the primitives of the model are such that the solution to this problem exists and is unique (or can be arbitrarily chosen from a set of candidate solutions; this assumption does not materially affect the points we raise in our discussion). Next, define
(4) $U_{\lambda}$ $=\mathbb{E}\big{[}U(\phi_{\lambda},\theta)\,\big{|}\,m\big{]}$ $V_{\lambda}$ $=\mathbb{E}\big{[}V(\phi_{\lambda},\theta)\,\big{|}\,m\big{]}$
which measure the amount of value that a particular choice of weight $\lambda$ accords to the agent and advocate, respectively. Note that, from the platform’s perspective, the total expected value generated at a particular value of $\lambda$ is given by $P_{\lambda}=U_{\lambda}+V_{\lambda}$. We are now ready to define our central measures of attentional agency.
Definition 0.
Pull is defined as the amount of informational value a platform delivers the agent as a fraction of the total value it is technologically capable of delivering them. Formally, for a given $u,v,\lambda$, $\Gamma$, and $m$, a platform’s pull is
(5) $\textbf{Pull}=\frac{U_{\lambda}}{U_{1}}$
Definition 0.
Push is the amount of informational value a platform delivers the advocate as a fraction of the total value it is technologically capable of delivering to them. Formally, for a given $u,v,\lambda$, $\Gamma$, and $m$, a platform’s push is
(6) $\textbf{Push}=\frac{V_{\lambda}}{V_{0}}$
Push and Pull both range from 00 to $1$ and index the amount of attentional value that is being delivered to agents and advocates, respectively. According to these definition, it is always the case that increasing push requires that one (weakly) decrease pull. Note, however, that push and pull are not necessarily zero-sum. We analyze the conditions that determine the relative tradeoffs between push and pull fully in the next section.
Another important point is that our framework defines push and pull for a single agent, advocate, and platform interaction. Across many interactions, each platform produces a distribution of Pull and Push values which could be analyzed in terms of their mean, variance, group-level disparities, and other distributional properties.
3.2. Analysis and Discussion
In this section we highlight some of its most salient aspects of the above framework and definitions.
3.2.1. Fixing technology, preference alignment determines the feasible push-pull frontier.
First, note that $U_{\lambda}$ is a weakly decreasing function of $\lambda$ and $V_{\lambda}$ is a weakly increasing function of $\lambda$ [^2] This is a direct consequence of the envelope theorem. However, the rate at which these functions change depends upon the relative alignment of agent and advocate preferences. Taking the platform’s allocation technology and agent model as given, we can consider the frontier of feasible push and pull values as a function of these underlying preferences. The platform effectively selects a point on this frontier by setting $\lambda$.
In a “zero-sum” situation where $u=-v$, then the platform’s objective simplifies to $(2\lambda-1)\,\mathbb{E}[U(\phi,\theta)|m]$, and they must choose either to deliver the agent’s most-preferred allocation (if $\lambda>\frac{1}{2}$) or their least-preferred allocation (if $\lambda<\frac{1}{2}$). In such an interaction, it is not only the case that pull and push are themselves zero sum ($\textbf{Pull}+\textbf{Push}=1$), but also that they change sharply around a critical value of $\lambda$. In other words, when agents and advocate push and pull in opposite directions, the platform will only accord agency to one of them. Conversely, if preferences are fully coincidental ($u=v$), then the platform can achieve maximal pull and push simultaneously ($\textbf{Pull}=\textbf{Push}=1$) for any value of $\lambda$ because the same allocation is optimal for both agents.
A less intuitive case of simultaneously maximal pull and push occurs when the agent is indifferent over certain aspects of their allocation that the advocate has strict preferences over and vice-versa. When preferences are “orthogonal” in this way, the platform can also achieve maximal pull and push simultaneously ($\textbf{Pull}=\textbf{Push}=1$) for any interior value of $\lambda$. From a policy perspective, situations of this variety have been described are ideal targets for “nudge”-style interventions (Thaler and Sunstein, 2009), given that they can, in theory, achieve a planner’s objectives without decreasing people’s attentional agency.
In the more general case where preferences are partially aligned, the platform must choose how much weight they wish to place on agent versus advocate informational value, and, accordingly, how much agency each should be allocated. Note, however, that from the platform’s perspective, the total informational value they create $P_{\lambda}$ is a concave function of $\lambda$ [^3] This is another direct consequence of the envelope theorem. This perhaps helps explain why interior values of Pull and Push are so commonly observed in practice—there is usually some value to the platform of choosing an interior $\lambda$ and dividing attentional agency to some extent.
3.2.2. Model advances increase platform value, but also redistribute attentional agency
Advances in machine learning and artificial intelligence (e.g., new architectures, more parameters, or additional compute (Kaplan et al., 2020)) enter our framework in two ways. The first is through $m$, which represents the depth and specificity with which the platform is able to infer an agent’s type $\theta$ before constructing an informational allocation.
Critically, however, such changes enhance the platform’s ability to provide value to both agent and advocate. In the former case, accurate inferences about an agent’s type enables the platform to more successfully extrapolate their preferences to hypothetical allocations. Simultaneously, these inferences also enable the platform to facilitate more precise targeting by advocates. Indeed, recent work shows that frontier models are capable of greater persuasion precisely because they are better at “matching the language or content of a message to the psychological profile of its recipient” (Matz et al., 2024), increasing the stakes of strategic misalignment between agents and advocates.
The net effect of model improvements on attentional agency therefore depend on whether the innovations are more useful for predicting marginal differences in $u$ or $v$ across various allocations. Recall that the platform’s objective when selecting a allocation is to maximize the linear combination
(7) $\lambda\,\mathbb{E}[U(\phi,\theta)\,|\,m]+(1-\lambda)\,\mathbb{E}[V(\phi,\theta)\,|\,m]$
for some value of $\lambda$. If two types $\theta$ and $\theta^{\prime}$ have conflicting preferences over allocations, then learning to distinguish them will make it more clear, in each case, what the marginal value of choosing one allocation over the other will be. This, in turn, will tend to shift emphasis in the overall optimization toward the agent’s preferences. A parallel argument, however, is also true for targeting.
Although the net effect on pull and push will depend, in each case, on which of these effects prevails, one thing is clear: the stakes of informational agency only grow as models become more sophisticated.
3.2.3. AI enables attentional allocation to operate over more granular and varied information objects
Advances in artificial intelligence also enter our framework through the partition $\Gamma$, which defines the units over which a platform can deliver information.
Multiple features of modern generative foundational models (Bommasani et al., 2021) now enable AI systems to achieve unprecedented levels of informational granularity: (1) attentional transformers parse a wide variety of data sources into their constitutive tokens, meaning that systems can apprehended complex informational objects at the scale of individual symbols (Vaswani et al., 2017); (2) ever-expanding context windows enable sequence-based models to capture long-range data dependencies (Press et al., 2021; Su et al., 2024); (3) multi-modality is in the process of extending the informational purview of AI systems to text, image, video, audio, and other formats (Yin et al., 2023); and (4) generative capabilities enable systems to flexibly recombine these informational atoms into arbitrary new structures that do not necessarily resemble the sources they were initially drawn from. Although it has always been possible for platforms to disassemble informational objects in a literal sense, foundation models enable a categorically greater degree of sophistication when extrapolating an agent’s preferences and intentions to these smaller units of information.
Within our framework, the cumulative effect of these innovations is to refine the platform’s informational domain into a finer partition $\Gamma^{\prime}\subset\Gamma$.  [^4] Specifically, if for every $\gamma^{\prime}\in\Gamma^{\prime}$ there exists a $\gamma\in\Gamma$ such that $\gamma^{\prime}\subseteq\gamma$ in the sense of weak subsetting, with at least one subset relationship strict. The additional degrees of freedom this opens up expand the space of possible allocations and, consequently, enable the platform to deliver more value overall. Formally, $\Phi(\Gamma)\subset\Phi(\Gamma^{\prime})$ so that ${P^{\prime}}_{\lambda}\geq P_{\lambda}$ for any given $\lambda$. As in the case of model development, the ultimate impact of such a change on welfare and informational agency will depend on the relative alignment of preferences and other primitives of the model.
One clear take-away, however, is that highly capable multi-modal AI models which are fully aligned to agents (i.e., for which $\lambda=1$ so that $\textbf{Pull}=1$) increase their ability to extract informational valuable from a variety of sources without satisfying the implicit “attentional bargain” that underlies many aspects of the internet’s current economic model—branding, sponsored content, advertising, and other features that businesses employ to monetize their content. As (Wu, 2017) points out in his history of the attention economy, the basic strategy to “draw attention with apparently free stuff and then resell it” has long been a part of the advertising business. Technological innovations intermittently disrupt the status quo implementation of this strategy, as exemplified by the famous “Betamax case” over whether Videocassette recordings of television shows (which enabled people to re-watch them without embedded advertisements) constituted copyright infringement. [^5] Sony Corp. of America v. Universal City Studios, Inc., 1984.
In contrast to prior waves of creative destruction, however, the very generality of general artificial intelligence means that one technology stands to disrupt nearly every existing informational format simultaneously—a development precedented only, perhaps, by advent of digital computers. The advent of human or super-human level artificial intelligence may represent a final frontier in these debates: by effectively solving the attention problem, these technology forces us to confront the incentive problem head on. We return to the question of how these insights inform policy approaches to refactoring the landscape of attentional agency in Section 5.3, below.
4. Attentional Agency in Present and Future Platforms
In the first part of this section, we describe how our framework maps on to five existing categories of internet platform: search, content libraries, social media, matching, and marketplaces. In the second part of this section, we apply our framework to consider the impact that advances in artificial intelligence are having on these platforms.
Search Content Social Matching Marketplace
Attentional Act Enter search query Browse library Refresh feed View potential matches Enter product query
Advocates Advertisers, SEO Advertisers, platforms Influencers, advertisers Promoted users Sellers, advertisers
Type
($\theta$)
Demographics, location, context Preferences over genres, formats, artists Social interests, browsing patterns Desired match, dating goals Shopping needs, market trends
Informational Objects
($X$, $\Gamma$)
Indexed websites, documents Books, music, movies, podcasts, shows Posts, stories, short-form video Dating profiles, job listings Product listings
Model
($m$)
Query, search history, IP address, cookies Viewing/listening history, ratings Behavioral analytics, engagement Preference matching Past purchases, viewing history, dynamic pricing
Table 1. How prominent categories of informational platforms map on to our formal framework.
Table 1 provides a high level overview of how various categories of platform map on to our framework. A side-by-side comparison reveals not only that many platforms are fundamentally oriented around attentional allocation, but that the boundaries separating these categories are becoming increasingly blurred over time. Search engines that may have once depended purely on a user’s entered query and a few contextual variables (e.g. location) now incorporate user history (Matthijs and Radlinski, 2011). Social media platforms often emphasize a feed (with minimal interaction options beyond “show me more content”) but may also support search. Online marketplaces sell a mixture of physical and digital goods, but now often offer recommendation features.
Below, we briefly highlight key characteristics of these five categories, then note how our model helps to clarify their conceptual similarities and differences. We then return to the topic of generative foundation models, which are beginning to blur some of the distinctions that have traditionally separated platform categories (Deldjoo et al., 2024).
4.1. Mapping Platform Categories Onto the Framework
Search Engine: Search engines rank web pages based on user queries (Varian, 2016; Resnick and Varian, 1997). Agents engaged in search generally want to specify their query as briefly as possible, meaning that search engines are most efficient when they can anticipate people’s needs and desires. This tends to incentivize platforms to build up a deep understanding of preferences and intention through repeated interaction with individual users. Interactions with the platform (e.g., click-through rates) are also used to tailor future results delivered to the same or other users (Resnick and Varian, 1997).
A central challenge of general web search is that pages are heterogeneous and constantly evolving, meaning that one must continuously crawl, index, and calculate the relevance of a very large number of informational objects to maintain relevance. As PageRank proved early on, the network structure of the web itself can be used as a proxy for relevance. Increasing levels of algorithmic sophistication have enabled search providers to predict the general quality and specific relevance of web pages with greater degrees of sophistication over time.
Content Library: Content libraries provide access to large collections of media and generally enable users to both search for items they already know and browse recommendations to discover new items. Media preferences tend to correlate significantly across individuals, meaning that general popularity is often a strong proxy for relevance to individual users. Personalization can significantly enhance the value of recommendations by inferring people’s particular tastes (e.g., a love of horror films) (Gomez-Uribe and Hunt, 2015).
Social Media: Recommendation-driven content feeds, such as those found on social media platforms, present users with a curated stream of multimedia objects centered around the activity of other users. Many social media feeds orient the user’s attention around a single activity: requesting more content. Social media platforms also tend to primarily infer an agent’s interests and intentions implicitly (e.g., from dwell time), rather than explicitly (e.g, entering a query). Social media is also distinguished by the fact that most individuals who act as agents also act as advocates: people both want to “see and be seen.”
Matching: Preferences on matching platforms are highly specific and driven by a clear objective, such as finding a compatible partner or securing a desired job. Matching platforms are also characterized by the fact that many users act as both agents and advocates. For example, successfully matching on a dating app requires other users to view and like one’s profile. Many matching platforms allow advocates (premium users) to gain visibility by influencing the attention of other users.
Marketplace : Online marketplaces help people search for goods and services. User preferences on marketplaces are driven by a combination of immediate shopping needs, long-term consumption habits, and sensitivity to factors like price, quality, and seller reputation. Attention can have a sizable impact on consumer behavior (Gabaix and Laibson, 2006), which makes featured listing, personalized recommendations, promotional offers, and targeted advertising especially valuable to advocates.
Summary: Although matching platforms are relatively easy to identify (items being ranked are platform participants themselves), marketplaces can be characterized by the actual exchange of consumer goods, and general search engines are relatively clearly defined, both content libraries and social media often offer a mixture of search, top-n recommendation, and infinite scroll content feeds.
Nonetheless, when we view each of these platforms as performing the core operation of delivery users ranked lists of some kind of item (be that web pages, videos, posts, other users, or offerings of physical goods), any of them can lean more towards pushing or pulling, and this decision is likely to be made on the basis of what is most profitable for that platform (barring policy interventions, such as those we discuss below).
4.2. The Effect of Generative Foundation Models
Table 2 summarizes how AI search tools (powered by generative foundation models) compare to the five existing platform categories just reviewed. As discussed in Section 3, these systems operate at the level of individual tokens (Bommasani et al., 2021), which enables them to flexibly tailor content to the user’s specific preferences ($\theta$), as inferred from their interaction history and context. The capacity of modern AI systems to draw meaningful connections between distal information sources and modalities (see Section 3.2.3) strengthens complementarities between various data sources: scientific papers, law articles, encyclopedia entries, code, and other sources of knowledge that were previously mostly referenced independently can now be interrelated in new and more valuable ways.
AI Search / Chat / Assistant
Attentional Act Request for information via a natural language prompt
Advocates (As of yet undetermined on most platforms)
Type ($\theta$) User preferences, needs, and intentions
Informational Domain ($X$, $\Gamma)$ Sub-components of any other informational domain
Model ($m$) Generative foundation model, context window, specific prompt
Table 2. Applying the push/pull framework to AI search, chat, and assistant tools, as powered by modern generative foundation models.
5. Discussion
5.1. Advertising in the Technology Industry
A common refrain in early criticisms of the targeted advertising model was “if you are not paying for the product, you are the product.” Subsequent proposals, such as data as labor (Posner and Weyl, 2018) and data dignity (Lanier and Weyl, 2018), pursue a related critique: that under the current paradigm, users of free internet platforms are also effectively workers producing data. Advocates of the advertising model, on the other hand, argue that it promotes equity and access because free products extend the benefits of technological innovation to those at the bottom of the socio-economic ladder.
Our framework shows how key claims frequently made in these debates can be made concrete. Our measures quantify the amount of attentional agency these platforms accord to various groups of individuals. If someone is not paying for a costly service, their activity is being subsidized in some way, often by advocates. The degree to which this reduces their attentional agency, however, depends on the relative alignment of preferences, technological context, and other factors.
This creates an open challenge for balancing the welfare implications of attentional agency and the consumer surplus afforded by ad-funded internet platforms (Brynjolfsson et al., 2023). Below, we outline policy ideas aimed at making tangible progress on this challenge, especially in light of recent AI progress.
5.2. Attentional Agency and Alignment
Our measures of attentional agency also complement ongoing research into AI alignment, especially work that emphasizes the challenge of aligning systems to the expressed or revealed values of diverse groups of stakeholders, using frames such as pluralism in human values (Sorensen et al., 2024) and social choice (Conitzer et al., 2024). Platforms create value by extending attention to previously unconsidered informational objects, which necessarily entails entrusting them with a certain degree of control over one’s attention. For AI-powered search and other systems that primarily function in this capacity, pull measures alignment to an individual user’s intentions. Push, on the other hand, measures alignment to other objectives, such as public trust and safety.
However, both push and pull represent special cases of more general conceptions of agency that are beginning to arise as AI systems become progressively more capable of handling decisions beyond information retrieval. A primary benefit of autonomous systems—i.e., ones that can act without needing to be continuously monitored—is that they free up people’s attention for other uses. This, however, necessarily reduces the amount of oversight such systems receive, which means they require substantially more trust in the faithful execution of user intentions and observance of public trust and safety objectives.
The concepts of push and pull clarify a variety of issues related to alignment and safety. For instance, interventions that aim to block specific concepts or pieces of information (e.g., instructions to build a weapon), such as preference ranking optimization (Song et al., 2024), or chained “safety classifiers” (Kim et al., 2023), and more can be understood as seeking to fulfill the goals of an advocate—which might be a public body acting in the public interest. In other words, alignment and safety interventions that are not highly personalized are best understood as supporting pushes from an advocate, often a representation of the public’s expectation of trust and safety. This insight is useful because modeling such situations as conflicts between advocates allows us to bring economic reasoning and mechanism design to bear on the problem.
5.3. Policy Strategies
There have been many proposals to promote greater transparency and accountability online. Some of the most prominent examples include the transparency provisions in the EU’s Digital Services Act (DSA), EU’s General Data Protection Regulation (GDPR), and proposed EU Artificial Intelligence Act (Economic and Committee, 2021; gdp, 2016)  [^6] Reporting of metrics that we outline in this paper are not new. For instance, Article 26 of the DSA requires platforms to provide users with meaningful information about the main parameters used in targeted advertising and ways to modify these settings. Article 26 also calls for a ban on targeted advertising based on profiling using special categories of personal data identified in Article 9 of the GDPR. Additionally, Article 27 of the DSA requires platforms to clearly explain the main parameters that influence what information is suggested to users and provide options to modify these parameters. Furthermore, Article 28 requires the largest online platforms and search engines to provide users with at least one recommender system option that is not based on profiling of individuals by automated systems as defined in the GDPR (Tech Policy Press, 2022).
Below, we highlight three policy strategies naturally suggested by our framework. Although a full implementation would require additional effort to account for platform-specific interactions with the framework, each of these relatively simple proposals stand to improve the distribution of informational agency in a measurable way.
  1. (1)
    Transparency into attentional agency: Measure and provide information on the distributions of Pull and Push values, both in aggregate and by subgroups. These disclosures would reveal particularly acute instances of misalignment between advocate and agent value on platforms. Greater transparency would be especially helpful for ensure that identifiable populations (e.g., people of a particular sexual orientation, racial identity, or age category) are experiencing different levels of attentional agency on a platform.
  2. (2)
    Expanded consumer choice: A variety of platforms sell push in order to generate revenue. For those that do, users could be given the option to pay for a “push-free” experience (which would generalize the “ad-free” tier offered by some services). An additional benefit of such options is that the revealed preferences of users who pursue it could inform standards for acceptable $\lambda$ values.
  3. (3)
    Provably neutral algorithms: Platforms offer a version of their service independent of advocate-driven factors, but that may not maximize pull. An example would be the inclusion of an option to sort one’s social media feed chronologically and without promotion.
While paid push-free options, neutral algorithm defaults, or push/pull disclosures may seem like relatively modest interventions, they reflect a fundamentally different theory of platforms than conventional transparency approaches. Rather than focusing on the black-box mechanics of ranking and recommendation systems, they target the concrete attentional choices that platforms make on behalf of users and the structural incentives that shape those choices. Disclosure-focused policy would open the door to broader societal conversations about equitable attentional agency.
These ideas are not a substitute for existing algorithmic audits, privacy protections, content moderation standards, or other key planks of the online safety agenda. Attentional agency attacks the problem of manipulative platform design from a neglected angle, one grounded in a numerical accounting of how platform incentives distort the information that users encounter. As policymakers grapple with the expanding domains in which algorithmic systems shape individual and collective experience, the push/pull framework helps identify and measure the core concept of attentional agency.
5.4. Impacts of Generative Foundation Models
Throughout this paper, we have commented on how our framework clarifies the likely effects of models that enable platforms to better understand users and to parse more granular units of information from across an increasing variety of sources. Given that these developments stand to disrupt the existing economic arrangement of the internet, it is also important to highlight how the above policy recommendations will interact with the present trajectory of AI development.
First, it could be the case that businesses avoid incorporating generative models into many of the platforms described above, in which our above recommendations remain highly relevant. Another possibility is that the ad-based funding model reasserts itself; perhaps ad-based search engines successfully integrate generative models while relying primarily on ad-based revenue, not subscriptions. In this scenario, transparency into $\lambda$ could become even more important, as ads become simultaneously more persuasive and less obviously ads. Finally, it could be the case that the rise of generative models catalyzes a general switch to subscription-based access to information. Of note, many leading AI efforts have chosen to rely on subscription and direct purchasing of API access. In this case, firms could see a stronger incentive to offer subscriptions, as this would be reflected in their informational agency disclosures.
Revenue generated by users paying to consume information could then be used to fund content creators. Although there are concerns in the short-term about potentially calamitous effects of language models on industries like journalism (Grynbaum and Mac, 2023), there exists a hopeful future in which direct access to highly personalized and carefully ranked chunks of information is worth paying for and ultimately leads to more funding allocated to information creation and curation. In this case, content creators may also become less reliant themselves on advertising. Achieving this hopeful future is, of course, highly path dependent, and will be impacted by regulatory outcomes in the near future and the ability for various information-providing industries to bargain with platforms that mediate information flow.
5.5. The Political Economy of Social Media and AI
Recent policy debates have highlighted a variety of issues that arise when large organizations have the capacity to determine what people pay attention to at scale. As one example, concerns about the political independence of TikTok’s content and algorithmic recommendations spurred the United States to pass legislation banning any “website, desktop application, mobile application, or augmented or immersive technology application” that is “controlled by a foreign adversary” (tik, 2024).
Our framework informs to these debates by providing a quantitative tool that can be used to evaluate the impact that competing incentives of various stakeholders have on user agency. By urging platforms to disclose push and pull metrics, policymakers can assess the actual balance of attentional agency on the platform and make evidence-based decisions about appropriate responses to emerging concerns.
The TikTok ban is just one example of the growing desire for better evidence-based approaches to platform accountability and governance. As just discussed in Section 5.4, generative foundation models and other innovations are quickly expanding the capacities of platforms, raising the stakes of developing a more participatory role for users in technological development (Allen and Weyl, 2024). By demanding transparency into the attentional agency of users, policymakers and the public can work towards aligning platform incentives with democratic values and the genuine preferences of users.
6. Conclusion
The framework we propose provides a lens for reasoning about how current platforms distribute attention. Our methods also provide new insight into ways that future platforms, especially ones powered by highly capable forms of artificial intelligence, might come closer to achieving the original visions of thinkers who inspired the internet. This framework we propose is intended to be readily applicable to policy discussions—indeed, the operational definitions we propose are based on techniques and computations already in use by platform operators.